### v2

@@ 3.6 Connections to delayed-loss models
We now compare the physical solution to the system (1.1)–(1.2) with an alternative class of models which we refer to as delayed-loss models. These models introduce a time lag in the propagation of losses between connected entities and, in their general form, are described by:
$X_{v}^{\lambda}(t)$ $=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}(t,X_{u}^{\lambda}),\quad v\in V$ (3.4)
$\tau_{v}$ $\mathrel{\mathop{\ordinarycolon}}=\tau(X_{v}^{\lambda})\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}\,X_{v}^{\lambda}\leq 0\},$ (3.5)
@@
$t\mapsto\lambda_{uv}(t,x)$ $\text{ is non-decreasing,}\quad\lambda_{uv}(t)=0\quad\forall t\in(-\infty,\tau(x)),\quad\lambda_{uv}(t)\leq 1\quad\forall t\in[\tau(x),\infty)$
$\text{where}\quad\tau(x)$ $\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}x(t)\leq 0\},$
systems (3.4) and (3.5) admit a unique solution.
@@
    where $(\xi_{v})_{v\in V}$ are i.i.d. exponential random variables with parameter $\kappa$ and $\kappa>0$ is a constant. $\lambda_{uv}(\cdot,x)$ converges in distribution to $\mathbf{1}_{[\tau(x),\infty)}(\cdot)$ as $\kappa\to\infty$.
@@
Let $(\lambda_{uv}^{n})_{(u,v)\in E}$ be a sequence of random functions indexed by $E$ such that $\lambda_{uv}^{n}(\cdot,x)$ converges locally uniformly around Brownian paths almost surely to $\mathbf{1}_{[\tau(x),\infty)}(\cdot)$ in $\mathcal{D}$ as $n\to\infty$ in the sense that
$\lim_{n\to\infty}d_{M_{1}}(\lambda_{uv}^{n}(\cdot,x^{n}),\mathbf{1}_{[\tau(x^{n}),\infty)}(\cdot))=0\quad\text{a.s.},$ (3.6)
@@
$B$ is the standard Brownian motion, and $(l^{n})_{n\geq 1}$ is any fixed sequence of uniformly bounded non-decreasing stochastic paths, for any $(u,v)\in E$. Then $(G,X^{\lambda^{n}},D^{n})$ converges almost surely to $(G,X,D)$ in $\mathcal{G}_{*}[\mathcal{D}^{2}]$ as $n\to\infty$.
@@
For each $n\geq 1$, we point out that the unique solution to (3.4) and (3.5) can be obtained by an iteration procedure similar to that in the construction of minimal solutions: $(G,X^{\lambda^{n}},D^{n})=\lim_{N\to\infty}(G,X^{\lambda^{n},N},D^{n,N})$, where
$X_{v}^{\lambda^{n},0}(t)=x_{v}+B_{v}(t),\quad X_{v}^{\lambda^{n},N+1}(t)=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t,X_{u}^{\lambda^{n},N})$
As
$X_{v}^{\lambda^{n},N+1}(t)$ $=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t,X_{u}^{\lambda^{n},N})\geq x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{n,N}\}},$
$0=X_{v}(\tau_{v})=x_{v}+B_{v}(\tau_{v})\geq x_{v}+B_{v}(\tau_{v})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}(t,X_{u}^{\lambda})=X_{v}^{\lambda^{n}},$
$x_{v}+B_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}+\varepsilon_{\Delta}<0.$
by equation (3.6). Therefore,
$X_{v}^{\lambda^{n}}(t_{\Delta})$ $=x_{v}+B_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t_{\Delta},X_{u}^{\lambda^{n}})$
$\leq x_{v}+B_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u}^{n},\infty)}(t_{\Delta})+\varepsilon_{\Delta}$
$\leq x_{v}+B_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}+\varepsilon_{\Delta}<0,$
To get the convergence of the solution paths $X^{\lambda_{n}}$, we note that
$X_{v}^{\lambda^{n}}(t)$ $=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t,X_{u}^{\lambda^{n}})$
$X_{v}(t)$ $=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u},\infty)}(t).$
$\leq d_{M_{1}}\Big{(}\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(\cdot,X_{u}^{\lambda^{n}}),\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u}^{n},\infty)}(\cdot)\Big{)}+d_{M_{1}}\Big{(}\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u}^{n},\infty)}(\cdot),\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u},\infty)}(\cdot)\Big{)}$
$\quad\to 0\quad\text{as}\quad n\to\infty.$
Therefore, we see that $X_{v}^{\lambda_{n}}$ converges almost surely to $X_{v}$ in $\mathcal{D}$ as $n\to\infty$ by combining the above limit with Lemma 2.3 and Lemma A.5. ∎
Acknowledgment. We are very grateful to Rene Carmona, Daniel Lacker, Sergey Nadtochiy, Mykhaylo Shkolnikov, H. Mete Soner and Ludovic Tangpi for offering insightful comments and suggestions.
Appendix A Appendix
@@
The following theorem is an easy extension of [36, Theorem 12.4.1] to $D([-1,\infty))$, the proof of which we omit for simplicity.
Theorem A.2.
@@
As each $f_{n}$ takes value in $\{0,1\}$ and is non-decreasing, its limit $f$ must take value in $\{0,1\}$ and be non-decreasing by using the fact that the continuity points are dense and $f$ is right-continuous. For any continuity point $t$ of $f$, we know from Theorem A.2 that $f(t)=\liminf_{n\to\infty}f_{n}(t)=0$ if $t<\limsup_{n\to\infty}\tau_{n}$, and $f(t)=\limsup_{n\to\infty}f_{n}(t)=1$ if $t>\liminf_{n\to\infty}\tau_{n}$. Combining the two cases, we force that $\liminf_{n\to\infty}\tau_{n}=\limsup_{n\to\infty}\tau_{n}$ and thus $\tau\mathrel{\mathop{\ordinarycolon}}=\lim_{n\to\infty}\tau_{n}$ exists and also that $f(t)=\mathbf{1}_{[\tau,\infty)}(t)$ by the right-continuity of $f$. ∎
Corollary A.4.
@@
The following technical lemma, which essentially says that addition is continuous at pairs of paths which do not jump in opposite directions at the same time, is useful for us.
Lemma A.5.
Suppose $f_{n}\to f$ and $g_{n}\to g$ in $D([-1,\infty))$ and that
$(f(t)-f(t-))(g(t)-g(t-))\geq 0$
for all $t\in[-1,\infty)$. Then $f_{n}+g_{n}\to f+g$ in $D([-1,\infty))$. In particular, the pointwise addition map $(c,f,g)\mapsto c+f+g$ is continuous from $\mathbb{R}\times C([-1,\infty))\times D([-1,\infty))$ to $D([-1,\infty))$.
@@
Combine [36, Theorem 12.7.3] with the definition of convergence in $D([-1,\infty))$. ∎

### v3

@@ 3.6 Connections to delayed-loss models
As the model studied in this paper features instantaneous propagation of losses, a natural question is how the physical solution compares with the solution to an alternative class of models which introduce a time lag in the propagation of losses between connected entities. Many of the latter, in their general form, can be described by:
$X_{v}^{\lambda}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}(t,X_{u}^{\lambda}),\quad v\in V$ (3.5)
$\tau_{v}$ $\mathrel{\mathop{\ordinarycolon}}=\tau(X_{v}^{\lambda})\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}\,X_{v}^{\lambda}\leq 0\},$
@@
$t\mapsto\lambda_{uv}(t,x)\text{ is non-decreasing,}$
$\lambda_{uv}(t)=0\quad\forall t\in(-\infty,\tau(x)),\quad\lambda_{uv}(t)\leq 1\quad\forall t\in[\tau(x),\infty)$
 $\quad\quad\quad\text{ where }\tau(x)\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}x(t)\leq 0\},$
system (3.5) admits a unique solution.
@@
    where $(\xi_{v})_{v\in V}$ are i.i.d. exponential random variables with rate parameter $\kappa>0$. $\lambda_{uv}(\cdot,x)$ converges in distribution to $\mathbf{1}_{[\tau(x),\infty)}(\cdot)$ as $\kappa\to\infty$.
@@
Let $(\lambda_{uv}^{n})_{(u,v)\in E}$ be a sequence of random functions such that, for every $(u,v)\in E$, the processes $\lambda_{uv}^{n}(\cdot,x)$ converges locally uniformly, around paths satisfying the downward crossing property, almost surely to $\mathbf{1}_{[\tau(x),\infty)}(\cdot)$ in $\mathcal{D}$ as $n\to\infty$, in the following sense:
$\lim_{n\to\infty}d_{M_{1}}\big(\lambda_{uv}^{n}(\cdot,x^{n}),\mathbf{1}_{[\tau(x^{n}),\infty)}(\cdot)\big)=0\qquad\text{a.s.},$ (3.6)
@@
and $Z$ is a fixed process satisfying the downward crossing property in Assumption 1.8, and $(l^{n})_{n\geq 1}$ is any fixed sequence of uniformly bounded, nondecreasing stochastic paths. Then $(G,X^{\lambda^{n}},D^{n})$ converges almost surely to $(G,X,D)$ in $\mathcal{G}_{*}[\mathcal{D}^{2}]$ as $n\to\infty$.
The proof will be given in Appendix C.
4 Concluding Discussion
Our analysis of fragility and robustness in default cascade models on infinite graphs provides insights into systemic risk in financial networks. In particular, the results highlight how network topology and local interactions fundamentally shape the resilience of banking systems to shocks. Highly interconnected institutions can serve as amplifiers of contagion, triggering widespread default cascades. This underscores the importance of incorporating network-based assessments into regulatory stress-testing frameworks, moving beyond institution-level risk metrics to system-wide vulnerability analyses.
The model also reveals that even minimal initial shocks can propagate extensively when structural fragility is present. This finding supports the case for targeted preventive measures—such as capital or liquidity buffers—for institutions that play a critical role in the network. More broadly, the framework offers a rigorous foundation for identifying systemic vulnerabilities and guiding preemptive strategies aimed at minimizing the probability and severity of financial crises.
4.1 Extensions to models with major banks
Admittedly, real financial networks often contain a small number of major institutions—such as central banks or globally systemic financial institutions—that interact with a large fraction of market participants. Such entities fall outside the present framework of local interactions on locally finite graphs, since they effectively introduce vertices of infinite degree and thereby distort the underlying graph distance, and hence the geometry, of the network. Nevertheless, the framework can be adapted by separating the global influence of these major institutions from the remaining local interactions among smaller banks.
One possible way to model such global effects is through the common noise component $Z^{0}$ as introduced in item 3 of Lemma 3.5. However, this approach has the drawback that it does not capture feedback from the system: in particular, it cannot account for how defaults of minor banks affect the financial health of major institutions.
Below we illustrate another more systematic way to incorporate this separation within our model. Minor banks continue to interact among themselves as in the previous formulation, but each minor bank also interacts simultaneously with all major banks. Thus, from the perspective of the minor banks, the major banks act as additional global (weak) counterparties. When the number of minor banks tends to infinity, the effect of the minor-bank population on any fixed major bank becomes diminishing (and potentially mean-field like), and therefore the major banks’ interaction structure could differ from that of the minor banks, depending on the specific modeling assumptions.
[…]
$\tilde{E}\mathrel{\mathop{\ordinarycolon}}=E\cup\{(i,v),(v,i)\mathrel{\mathop{\ordinarycolon}}v\in V,i\in[N]\}\cup\{(i,j)\in[N]^{2}\mathrel{\mathop{\ordinarycolon}}i\neq j\}.$
The corresponding interacting particle system on $\tilde{G}$ is
$X_{v}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}}-\sum_{i=1}^{N}c_{iv}\mathbf{1}_{\{i\in D_{t}\}},\quad v\in V$ (4.1)
$X_{i}(t)$ $=x_{i}+Z_{i}(t)-\sum_{v\in V}c_{vi}\mathbf{1}_{\{v\in D_{t}\}}-\sum_{j=1}^{N}c_{ji}\mathbf{1}_{\{j\in D_{t}\}},\quad i\in[N]$
$D_{t}$ $=\{v\in\tilde{V}\mathrel{\mathop{\ordinarycolon}}\,\inf_{s\in[0,t]}X_{v}(s)\leq 0\},$
where it is necessary to impose summability conditions
$\max_{i\in[N]}\,\sum_{v\in V}\,(c_{vi}+c_{iv})<\infty.$
With a suitable refinement of the arguments developed in this chapter (in particular, the percolation argument should still be performed on the graph $G$ rather than $\tilde{G}$), it should be possible to extend the main results to the system (4.1). We leave this extension for future work.
4.2 Other future directions
We comment on several promising directions for future research that emerge from this work. A first avenue involves enriching the model with detailed balance sheet structures, moving beyond a scalar health variable to explicitly account for asset-liability compositions, claim seniority, collateralization, and regulatory capital buffers. This would allow for more granular estimation of creditor losses and more realistic modeling of contagion mechanisms.
Second, introducing dynamic network features—such as institutional entry and exit, multiple defaults per node, and adaptive exposure reallocation—would better capture the evolving nature of financial systems. Incorporating regulatory constraints, market frictions, and mechanisms for resolution and re-entry would further enhance the framework’s ability to reflect real-world institutional behavior.
Third, embedding strategic interaction into network formation would allow institutions to optimize exposures under risk-return trade-offs, subject to systemic constraints. A game-theoretic extension could capture how such strategic decisions interact and how macro-prudential tools influence network architecture and collective outcomes.
Finally, empirical calibration using data on interbank exposures and historical default events is essential for improving the model’s practical relevance. Estimating unobserved network structures, calibrating parameters to match crisis dynamics, and validating the model against real contagion episodes would enhance its usefulness for central banks and financial regulators engaged in stress testing and systemic risk monitoring.
Appendix A The space $\mathcal{D}\mathrel{\mathop{\ordinarycolon}}=D([-1,\infty))$ and the $M_{1}$-topology
A.1 The space $D([a,b])$
A.2 The space $\mathcal{D}\mathrel{\mathop{\ordinarycolon}}=D([-1,\infty))$
Lemma A.1.
A.3 Towards compactness in $\mathcal{D}$
The next result provides a sufficient condition for compactness of subsets of $D([-1,\infty))$ consisting of monotone and uniformly bounded paths. This compactness criterion will be useful in proving tightness of empirical processes arising from our particle system.
Theorem A.2.
$\mathcal{M}((M_{m})_{m\geq 1})\mathrel{\mathop{\ordinarycolon}}$$\{$ $f\in D([-1,\infty))\mathrel{\mathop{\ordinarycolon}}\,f=0\,\,\text{on }[-1,0),$
$\text{$f$ is monotone and }\sup_{[-1,m]}|f|\leq M_{m},\quad\forall m\geq 1\}.$
In particular, for any $M\in(0,\infty)$, the set
has compact closure in the $M_{1}$ topology.
@@ A.3 Towards compactness in $\mathcal{D}$
which is monotone and remains constant on [1,0)[-1,0) and on $(m,m+1]$. By Theorem [34, Theorem 12.12.2], there exists $f_{\infty}^{m}\in D([-1,m+1])$ such that $f_{k}^{m}$ converges to $f_{\infty}^{m}$ in D([1,m+1])D([-1,m+1]) as $k\to\infty$ along some subsequence, which is potentially a further subsequence of the subsequence corresponding to m1m-1. In particular, by [34, Corollary 12.9.1], we can find a continuity point $t_{m}\in(m-1,m]$ of $f_{\infty}^{m}$ such that $f_{k}|_{[-1,t_{m}]}=f_{k}^{m}|_{[-1,t_{m}]}$ converges to $f_{\infty}^{m}|_{[-1,t_{m}]}$ in $D([-1,t_{m}])$ as $k\to\infty$ along that subsequence (note also that $t_{m}\uparrow\infty$ by construction). Combining the diagonal argument, Theorem B.2 and the right-continuity of the limit function, we can find an $f_{\infty}\in D([-1,\infty))$ such that $f_{k}$ converges to $f_{\infty}$ in $D([-1,\infty))$ as $k\to\infty$ along some subsequence. ∎
Appendix B Auxiliary Technical Results
@@
The following theorem is an easy extension of [34, Theorem 12.4.1] to $D([-1,\infty))$.
Theorem B.2.
@@
As each $f_{n}$ takes value in $\{0,1\}$ and is non-decreasing, its limit $f$ must take value in $\{0,1\}$ and be non-decreasing by using the fact that the continuity points are dense and $f$ is right-continuous. For any continuity point $t$ of $f$, we know from Theorem B.2 that $f(t)=\liminf_{n\to\infty}f_{n}(t)=0$ if $t<\limsup_{n\to\infty}\tau_{n}$, and $f(t)=\limsup_{n\to\infty}f_{n}(t)=1$ if $t>\liminf_{n\to\infty}\tau_{n}$. Combining the two cases, we force that $\liminf_{n\to\infty}\tau_{n}=\limsup_{n\to\infty}\tau_{n}$ and thus $\tau\mathrel{\mathop{\ordinarycolon}}=\lim_{n\to\infty}\tau_{n}$ exists and also that $f(t)=\mathbf{1}_{[\tau,\infty)}(t)$ by the right-continuity of $f$. ∎
Corollary B.4.
@@
The following two Lemmas characterize the “uniform convergence” entailed by convergence in the $M_{1}$ topology.
Lemma B.5.
Suppose that the sequence $(f_{n})_{n\geq 0}$ converges to $f$ in $D([-1,\infty))$ in the $M_{1}$ topology, and $t_{n}\to t\geq 0$. Then
$f(t-)\wedge f(t)\leq\liminf_{n\to\infty}f_{n}(t_{n}-)\wedge f_{n}(t_{n})\leq\limsup_{n\to\infty}f_{n}(t_{n}-)\vee f_{n}(t_{n})\leq f(t-)\vee f(t).$
@@
For any $\delta>0$, we can take $\xi_{1}\in(t-\delta,t)$ and $\xi_{2}\in(t,t+\delta)$ that are continuity points of $f$. Noting the definition of $w_{s}(f,t,\delta)$ as in [34, Equation (12.4.4)], we get
$f_{n}(\xi_{1})\wedge f_{n}(\xi_{2})-w_{s}(f_{n},t,\delta)\leq f_{n}(t_{n}-),f_{n}(t_{n})\leq f_{n}(\xi_{1})\vee f_{n}(\xi_{2})+w_{s}(f_{n},t,\delta).$
Taking $n\to\infty$ and then $\delta\downarrow 0$, we get the desired result by [34, Theorem 12.5.1 (iv)]. ∎
Lemma B.6.
Suppose that the sequence $(f_{n})_{n\geq 0}$ converges to $f$ in $D([-1,\infty))$ in the $M_{1}$ topology, and that $a^{\prime},b^{\prime}]\subset(a,b)\subset[-1,\infty)$. Then
$\inf_{t\in[a,b]}f(t)\leq\liminf_{n\to\infty}\inf_{t\in[a^{\prime},b^{\prime}]}f_{n}(t)\leq\limsup_{n\to\infty}\sup_{t\in[a^{\prime},b^{\prime}]}f_{n}(t)\leq\sup_{t\in[a,b]}f(t).$
Proof.
Take $\delta>0$ sufficiently small so that $a^{\prime}-\delta,b^{\prime}+\delta]\subset[a,b]$. We can find a partition $t_{0}<t_{1}<...<t_{M}$ consisting of continuity points of $f$ such that $t_{0}\in[a^{\prime}-\delta,a^{\prime}]$, $t_{M}\in[b^{\prime},b^{\prime}+\delta]$, and $t_{i+1}-t_{i}\leq\delta$. For any $t\in[a^{\prime},b^{\prime}]$, let $i$ be such that $t\in[t_{i},t_{i+1}]$. Noting the definition of $w_{s}(f,\delta)$ as in [34, Equation (12.5.1)] (we can first restrict each $f_{n}$ to the compact interval $1,b+\delta]$, to avoid taking supremum of $w_{s}(f_{n},t,\delta)$ over the non-compact interval $1,\infty)$), we get
$\min\{f_{n}(t_{i}),f_{n}(t_{i+1})\}-w_{s}(f_{n},\delta)\leq f_{n}(t)\leq\max\{f_{n}(t_{i}),f_{n}(t_{i+1})\}+w_{s}(f_{n},\delta).$
As a result,
$\min_{0\leq i\leq M}f_{n}(t_{i})-w_{s}(f_{n},\delta)\leq\inf_{t\in[a^{\prime},b^{\prime}]}f_{n}(t)\leq\sup_{t\in[a^{\prime},b^{\prime}]}f_{n}(t)\leq\max_{0\leq i\leq M}f_{n}(t_{i})+w_{s}(f_{n},\delta).$
Taking $n\to\infty$ and then $\delta\downarrow 0$, we get the desired result by combining [34, Theorem 12.5.1 (iv)] and Theorem B.2. ∎
Appendix C Remaining Technical Proofs
Proof of Proposition 3.16.
$\underline{F}\mathrel{\mathop{\ordinarycolon}}=\lim_{n\to\infty}F^{n}$
exists in the sense of weak convergence of CDFs and gives the minimal fixed point of $\Psi$. To interpret this iteration probabilistically, define the truncated systems $G_{n}=G|_{[-n,n]}$ with $V_{n}\mathrel{\mathop{\ordinarycolon}}=\{i\in\mathbb{Z}\mathrel{\mathop{\ordinarycolon}}-n\leq i\leq n\}$ and $E_{n}\mathrel{\mathop{\ordinarycolon}}=\{(i,i+1)\mathrel{\mathop{\ordinarycolon}}-n\leq i<n\}$. Let the roots of $G_{n}$ and $G$ be the vertex 0, and define $(c^{n},x^{n},Z^{n})$ on $G_{n}$ by the corresponding restriction of $(c,x,Z)$. Apparently, $(G_{n},c^{n},x^{n},Z^{n})$ converges almost surely to $(G,c,x,Z)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}]$ as $n\to\infty$, which, in particular, implies $\mathcal{L}(\tau_{0}^{n})\to\mathcal{L}(\tau_{0})$ as $n\to\infty$ by Theorem 1.11 and Corollary B.3. As the vertex n-n has no in-neighbors in $G_{n}$, we see that
$\mathbb{P}[\tau_{-n}^{n}\leq t]=\mathbb{P}\Big[\inf_{s\in[0,t]}(x_{-n}+Z_{-n}(s))\leq 0\Big],$
that is, $F_{\tau_{-n}^{n}}=\Psi[F^{0}]$. Moreover, as
$X_{i+1}^{n}(t)=x_{i+1}+Z_{i+1}(t)-c_{i,i+1}\mathbf{1}_{\{\tau_{i}^{n}\leq t\}}$
and that $\tau_{i}^{n}$ is independent of $(c_{i,i+1},x_{i+1},Z_{i+1})$, the latter of which has the same distribution as $(c_{-1,0},x_{0},Z_{0})$, it holds that $F_{\tau_{i+1}^{n}}=\Psi[F_{\tau_{i}^{n}}]$ for $i\geq-n$. As a result, we obtain the relation $F_{\tau_{0}^{n}}=\Psi^{n}[F_{\tau_{-n}^{n}}]=F^{n+1}$, and the claim is proved. ∎